OpenStudy (anonymous):

How do i graph the polynomial function:f(x) = -x^(3) - 2

OpenStudy (anonymous):

Find three sets that solve the equation

OpenStudy (mertsj):

Find the basic shape of a cubic. Move it down two units and reflect it about the x axis since the coefficient of the x^3 term is negative.

OpenStudy (anonymous):

Could one of you do a step by step for me?

OpenStudy (anonymous):

I'm having hard time remembering but I'll see what I can do, first replace the x with a number lets say 1 so the equation would look like this f(x)=-1^(3)-2..... f(x)=-1-2.......f(x)=3, so the set is (1,3), keep doing this till you get three sets then check. P.S. To those who might answer after me if I did something wrong tell me.

OpenStudy (anonymous):

so i can pick any number to use?

OpenStudy (anonymous):

from what I remember yes just make sure you don't screw up when you solve the equation that could throw you off.

OpenStudy (anonymous):

give me an example and I'll see if your doing it right

OpenStudy (anonymous):

f(x)=-2^(3)-2 =-8-2 =-10 (2,-10)